# Structural Entropy Guided Unsupervised Graph Out-Of-Distribution Detection

### v1

@@ Introduction
Graph Contrastive Learning. As an effective graph self-supervised learning paradigm (Liu et al. 2021, 2022), GCL has achieved great success on unsupervised graph representation learning (Velickovic et al. 2020; Sun et al. 2020a; Hassani and Khasahmadi 2020; You et al. 2020; Zhu et al. 2021; Qiu et al. 2020; Zheng et al. 2022a, b; Ding et al. 2022). Typically, GCL methods involve generating diverse graph views through data augmentation techniques and optimizing the mutual agreement between these views to enhance the representation of samples with similar semantic semantics (You et al. 2020; Zhu et al. 2021; Hassani and Khasahmadi 2020; Zheng et al. 2022b; Ding et al. 2022). However, methods that perform augmentation on graph structures may inadvertently introduce undesired OOD samples within ID data, and views that enhance graph features often suffer from containing redundant information. To effectively capture the essential structure of original graphs, this paper introduces a GCL framework guided by structural entropy, innovatively incorporating triplet views and multi-grained contrast.
@@
Before formulating the research problem, we first provide some necessary notations. Let $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. The node features are represented by the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d_{f}}$, where $n=|\mathcal{V}|$ is the number of nodes and $d_{f}$ is the feature dimension. The structure information can also be described by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, so a graph can be alternatively represented by $G=(\mathbf{A},\mathbf{X})$. Notation with its description can be found in Appendix A.
@@
The detailed proof of Theorem 6 is shown in Appendix B.1. Here, the mutual information between $G$ and $G^{\ast}$ can be rewritten as:
@@
With the essential view eliminating redundant information, we also theoretically prove that our SEGO effectively captures the maximum mutual information between the representations obtained from the anchor view and labels by minimizing the contrastive loss in Eq.3. The InfoMax principle has been widely applied in representation learning literature (Bachman, Hjelm, and Buchwalter 2019; Poole et al. 2019; Tschannen et al. 2020; Yuan et al. 2024a). MI quantifies the amount of information obtained about one random variable by observing the other random variable. We first introduce two lemmas, which are proven in Appendix B.2.
@@
Theorem 2 (See proof in Appendix B.3) reveals that our anchor view empowers the encoder to acquire enhanced representations through contrastive learning and preserve more information associated with the ground-truth labels, which will boost the performance of downstream tasks.
@@
$\mathcal{H}^{(k)}({G})=\min_{\forall T:\mathrm{Height}(T)=k}\{\mathcal{H}^{T}({G})\}.$ (10)
The total process of generation of a coding tree with fixed height $k$ can be divided into two steps: 1) construction of the full-height binary coding tree and 2) compression of the binary coding tree to height $k$. Given root node $v_{r}$ of the coding tree $T$, all original nodes in graph ${G}=(\mathcal{V},\mathcal{E})$ are treated as leaf nodes. Correspondingly, based on SEP (Wu et al. 2022a), we design two efficient operators, $\mathbf{MERGE}$ and $\mathbf{DROP}$, to construct a coding tree $T$ with minimum structural entropy. This process is detailed in Algorithm 1 in Appendix C. This tree anchor view $T=\arg\min{\mathcal{H}^{T}({G_{b}})}$ effectively removes redundant information from graphs while preserving distinctive essential structural information, enabling the capture of distinct graph patterns between ID and OOD samples.
@@
In this section, we empirically evaluate the effectiveness of the proposed SEGO.  [^1] The code of SEGO is available at https://github.com/name-is-what/SEGO. In particular, the experiments are unfolded by answering the following research questions:
@@
Datasets. For OOD detection, we employ 10 pairs of datasets from two mainstream graph data benchmarks (i.e., TUDataset (Morris et al. 2020) and OGB (Hu et al. 2020)) following GOOD-D (Liu et al. 2023). We also conduct experiments on anomaly detection settings, where 5 datasets from TUDataset (Morris et al. 2020) are used for evaluation, where the samples in minority class or real anomalous class are viewed as anomalies, while the rest are viewed as normal data. Further details are shown in Appendix D.1.
Baselines. We compare SEGO with 14 competing baseline methods, including 6 GCL (Sun et al. 2020a; You et al. 2020; Liu et al. 2023) based methods, 6 graph kernel based methods (Vishwanathan et al. 2010; Shervashidze et al. 2011), and 2 end-to-end graph anomaly detection methods (Zhao and Akoglu 2021; Ma et al. 2022). More details about implementation are provided in Appendix D.2.
@@
To answer RQ1, we compare our proposed methods with 14 competing methods in OOD detection tasks. The AUC results are reported in Table 1. From the comparison results, we observe that SEGO achieves superior performance improvements over the baselines. Specifically, SEGO achieves the best performance on 9 out of 10 dataset pairs and ranks first on average among all baselines with an average rank (A.R.) of 1.1. Additionally, SEGO outperforms all the compared methods in terms of average AUC with a score of 84.46, which is 3.7% higher than the second-best method GOOD-D (Liu et al. 2023). Notably, on the FreeSolv/ToxCast dataset pair, SEGO surpasses the best competitor by 10.8%. While SEGO may not have achieved the absolute best results on the IMDB-M/IMDB-B datasets, its performance is nearly optimal. This is likely because the current coding tree is an approximation that doesn’t fully eliminate all irrelevant redundant information for the high connectivity and edge density of social networks.
@@
To investigate if SEGO can generalize to the anomaly detection setting (Zhao and Akoglu 2021; Ma et al. 2022), we conduct experiments on 5 datasets following the benchmark in GLocalKD and GOOD-D (Ma et al. 2022; Liu et al. 2023), where only normal data are used for model training. The results are shown in Table $2$More experiments on other 5 datasets for anomaly detection can be found in Appendix D.3. From the results, we find that SEGO shows significant performance improvements compared to other baseline methods. Capturing common patterns in the anomaly detection setting is crucial, which is directly reflected in the performance. Thus, we can conclude that SEGO indeed has a strong capability to learn the essential information of normal graph data.
@@
The Height $k$ of Graph’s Natural Hierarchy. In the experimental setup, the height $k$ of the coding tree is consistently set to 5, in alignment with the GNN encoder. Here, we delve deeper into the effect of the height $k$ on the graph’s natural hierarchy. The specific performance of SEGO under each height $k$, ranging from 2 to 5, on OOD detection is shown in Fig. A.5. We can observe that the optimal height $k$ with the highest accuracy varies among datasets. We also conducted experiments on the impact of the coding tree height on anomaly detection datasets, as detailed in Appendix D.4.

### v2

@@ Introduction
Graph Contrastive Learning. As an effective graph self-supervised learning paradigm (Liu et al. 2021, 2022), GCL has achieved great success on unsupervised graph representation learning (Velickovic et al. 2020; Sun et al. 2020; Hassani and Khasahmadi 2020; You et al. 2020; Zhu et al. 2021; Qiu et al. 2020; Zheng et al. 2022a, b; Ding et al. 2022). Typically, GCL methods involve generating diverse graph views through data augmentation techniques and optimizing the mutual agreement between these views to enhance the representation of samples with similar semantic semantics (You et al. 2020; Zhu et al. 2021; Hassani and Khasahmadi 2020; Zheng et al. 2022b; Ding et al. 2022). However, methods that perform augmentation on graph structures may inadvertently introduce undesired OOD samples within ID data, and views that enhance graph features often suffer from containing redundant information. To effectively capture the essential structure of original graphs, this paper introduces a GCL framework guided by structural entropy, innovatively incorporating triplet views and multi-grained contrast.
@@
Before formulating the research problem, we first provide some necessary notations. Let $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}$ is the set of edges. The node features are represented by the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d_{f}}$, where $n=|\mathcal{V}|$ is the number of nodes and $d_{f}$ is the feature dimension. The structure information can also be described by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, so a graph can be alternatively represented by $G=(\mathbf{A},\mathbf{X})$.
@@
Here, the mutual information between $G$ and $G^{\ast}$ can be rewritten as:
@@
With the essential view eliminating redundant information, we also theoretically prove that our SEGO effectively captures the maximum mutual information between the representations obtained from the anchor view and labels by minimizing the contrastive loss in Eq.3. The InfoMax principle has been widely applied in representation learning literature (Bachman, Hjelm, and Buchwalter 2019; Poole et al. 2019; Tschannen et al. 2020; Yuan et al. 2024a). MI quantifies the amount of information obtained about one random variable by observing the other random variable. We first introduce two lemmas below.
@@
Theorem 2 reveals that our anchor view empowers the encoder to acquire enhanced representations through contrastive learning and preserve more information associated with the ground-truth labels, which will boost the performance of downstream tasks.
@@
$\mathcal{H}^{(k)}({G})=\min_{\forall T:\mathrm{Height}(T)=k}\{\mathcal{H}^{T}({G})\}.\vspace{-1.0mm}$ (10)
The total process of generation of a coding tree with fixed height $k$ can be divided into two steps: 1) construction of the full-height binary coding tree and 2) compression of the binary coding tree to height $k$. Given root node $v_{r}$ of the coding tree $T$, all original nodes in graph ${G}=(\mathcal{V},\mathcal{E})$ are treated as leaf nodes. Correspondingly, based on SEP (Wu et al. 2022a), we design two efficient operators, $\mathbf{MERGE}$ and $\mathbf{DROP}$, to construct a coding tree $T$ with minimum structural entropy. This tree anchor view $T=\arg\min{\mathcal{H}^{T}({G_{b}})}$ effectively removes redundant information from graphs while preserving distinctive essential structural information, enabling the capture of distinct graph patterns between ID and OOD samples.
@@
In this section, we empirically evaluate the effectiveness of the proposed SEGO. In particular, the experiments are unfolded by answering the following research questions:
@@
Datasets. For OOD detection, we employ 10 pairs of datasets from two mainstream graph data benchmarks (i.e., TUDataset (Morris et al. 2020) and OGB (Hu et al. 2020)) following GOOD-D (Liu et al. 2023). We also conduct experiments on anomaly detection settings, where 5 datasets from TUDataset (Morris et al. 2020) are used for evaluation, where the samples in minority class or real anomalous class are viewed as anomalies, while the rest are as normal data.
Baselines. We compare SEGO with 14 competing baseline methods, including 6 GCL (Sun et al. 2020; You et al. 2020; Liu et al. 2023) based methods, 6 graph kernel based methods (Vishwanathan et al. 2010; Shervashidze et al. 2011), and 2 end-to-end graph anomaly detection methods (Zhao and Akoglu 2021; Ma et al. 2022).
@@
To answer RQ1, we compare our proposed methods with 14 competing methods in OOD detection tasks. The AUC results are reported in Table 1. From the comparison results, we observe that SEGO achieves superior performance improvements over the baselines. Specifically, SEGO achieves the best performance on 9 out of 10 dataset pairs and ranks first on average among all baselines with an average rank (A.R.) of 1.1. Additionally, SEGO outperforms all the compared methods in terms of average AUC with a score of 84.46, which is 3.7% higher than the second-best method GOOD-D (Liu et al. 2023). Notably, on the FreeSolv/ToxCast dataset pair, SEGO surpasses the best competitor by 10.8%. Although SEGO nearly achieves optimal results on the IMDB-M/IMDB-B datasets, it falls short likely because its coding tree only approximates and doesn’t fully remove redundant information. Additionally, the high connectivity and edge density of social network datasets introduce more redundancy, making it harder to capture essential information. These results underscore the superiority of SEGO in OOD detection tasks, demonstrating its ability to capture essential information across different granular levels.
@@
To investigate if SEGO can generalize to the anomaly detection setting (Zhao and Akoglu 2021; Ma et al. 2022), we conduct experiments on 5 datasets following the benchmark in GLocalKD and GOOD-D (Ma et al. 2022; Liu et al. 2023), where only normal data are used for model training. The results are shown in Table 2. From the results, we find that SEGO shows significant performance improvements compared to other baseline methods. Capturing common patterns in the anomaly detection setting is crucial, which is directly reflected in the performance. Thus, we can conclude that SEGO indeed has a strong capability to learn the essential information of normal graph data.
@@
The Height $k$ of Graph’s Natural Hierarchy. In the experimental setup, the height $k$ of the coding tree is consistently set to 5, in alignment with the GNN encoder. Here, we delve deeper into the effect of the height $k$ on the graph’s natural hierarchy. The specific performance of SEGO under each height $k$, ranging from 2 to 5, on OOD detection is shown in Fig. 5. We can observe that the optimal height $k$ with the highest accuracy varies among datasets.